i have 30 coins in my pocket, some dimes, twice as many nickels, and the rest quareters. the total value is $5.30. how many dimes do i have?
Answer by mananth(16949)

dimes=x
nickels=2x
quarters = 30-(x+2x)
=30-3x.
..
10x+5*2x+25(30-3x)=530
10x+10x+750-75x=530
-55x=530-750
-55x=-220
x=4 dimes
nickels = 2*4 =8
quarters = 30-3*4= 18
...
4*10+8*5+18*25=530